Suppose that market demand and supply function of product x are as follows DD=1000+40P
SS=600+4P what is the equilibrium price and equilibrium quantity of x
1
Expert's answer
2020-12-14T05:10:23-0500

The equilibrium price and equilibrium quantity of x are:

DD = SS,

"1000 - 40P = 600 + 4P,"

P = 9.1,

x = 600 + 4×9.1 = 636.36 units.